Solve the linear equations 2x+4y=10 and 3x+6y=15.
Answer by richwmiller(17219)

Both equations represent the same line so there are infinite solutions.
Divide this one by 2: 2x+4y=10
Divide this one by 3: 3x+6y=15
and you get x+2y=5 for both